How much of an alloy that is 20% copper should be mixed with 400 ounces of an alloy that is 80% copper in order to get an alloy that is 50% copper?
Found 2 solutions by addingup, adamhen894:
Answer by addingup(3677)

0.2x+0.8(400) = 0.5(x+400)
0.2x+320 = 0.5x+200)
-0.3x = -120
x = 400 You need 400 oz of the 20%

Answer by adamhen894(15)

x ounces of 20% copper
mix with
400 ounces of 80% copper
to get
(x+400) ounces of 50% copper
set up the equation, solve for x
0.2x+0.8(400)=0.5(x+400)
x = 400 ounces
if you input x in the equation, it equals on both side.
if you have any alternative solution, i would love to hear.
adamchen894@gmail.com